the width of a rectangle is 5 feet less than its length. Find its width if the perimeter is 54 feet.

Answer by stanbon(75887)

the width of a rectangle is 5 feet less than its length. Find its width if the perimeter is 54 feet.
-----------
length = L
width = L-5
------
Perimeter = 2(L+W)
54 = 2(L + L-5)
27 = 2L - 5
2L = 32
L = length = 16 ft.
W = width = L-5 = 11 ft
=================
Cheers,
Stan H.
====================
--------------